OpenStudy (anonymous):

A car dealership pays you 4% commission on your first $8000 in car sales and 8% commission on sales over $8000. If you made $9,000 in car sales, how much is your commission?

OpenStudy (anonymous):

Any body know how to do this.

OpenStudy (whpalmer4):

Yes :-)

OpenStudy (anonymous):

(0.04)(8000) + (0.08)(9000 - 8000)

OpenStudy (whpalmer4):

Set up equations to show up much you make. For sales <= $8000, you get 4%, and for sales > $8000 you get 8%.

OpenStudy (anonymous):

That becomes: (0.04)(8000) + (0.08)(1000) 320 + 80 = 400

OpenStudy (anonymous):

Do I turn he percent into a decimal

OpenStudy (whpalmer4):

Yes. x% is the same as x/100. So 4% = 4/100 = 0.04.

OpenStudy (anonymous):

In my calculations, the 0.04 is 4% and 0.08 is 8%

OpenStudy (anonymous):

"9000 - 8000" represents the amount over 8000 which is 1000

OpenStudy (anonymous):

so whats that supposed to mean. I'm confused

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

So subtract 8000 from 9000 and get 1000 then what do I do

OpenStudy (anonymous):

Just look at my first 2 posts. There are 1000 in sales over and above 8000 because there were 9000 in sales total.